OpenStudy (anonymous):

A sample of gas has a pressure of 3.2 atmospheres at a temperature of 301 Kelvin. What would be the pressure of the same sample of gas when the temperature is decreased to 285 Kelvin?

OpenStudy (abmon98):

Use P1/P2=T1/T2 3.2/P2=301/285